OpenStudy (errrwut):

Need help understanding how an answer was achieved: (my worked out answer is inside) Particle physicists use particle track detectors to determine the lifetime of short-lived particles. A muon has a mean lifetime of 2.2 micro/seconds and makes a track 9.5 cm long before decaying into an electron and two neutrinos. What was the speed of the muon?

OpenStudy (errrwut):

\[T_0 = 2.2 µs\] \[L = 9.5cm\] Using, \[\beta = \frac{L*\sqrt{1-\beta^2}}{c*T_0}\] Solving for \[\beta\] \[\beta=\frac{L}{\sqrt{L^2+c^2*T^2}}\] You can neglect c for the mean time? I get \[\beta = 0.974218\] However, the solution is supposedly \[\beta = 1.4*10^{-4}\]

OpenStudy (errrwut):

I also posted more information here: http://openstudy.com/updates/5807fa6ee4b05a233fe7f33c